# A Differentiable Model of Nucleic Acid Dynamics

### v1

@@ 2.3 Mechanical Optimization
Figure 4: Gradient-based parameter optimization for mechanical properties in oxDNA. A. Our method for parameter optimization for complex target mechanical properties via DiffTRE and implicit derivatives. B. Parameter optimization for target values of persistence length ($L_{ps}$). I. A depiction of $L_{ps}$. II. Optimization of all oxDNA parameters to match a set target values of $L_{ps}$. III. The top three most important parameters ranked via the absolute value of the gradient. C. Parameter optimization for target values of moduli characterizing response to external force and torque. I. A depiction of a duplex under external force and torque. II-III. Optimization of all oxDNA parameters to match the twist-stretch coupling, torsional modulus, and effective stretching modulus, all within experimental error. Outliers in the twist-stretch coupling optimization at iterations 52 and 53 are omitted and replaced with a dashed line. See Figure S6 for corresponding sensitivity analyses.
@@
To solve these challenges, we extend DiffTRE to compute derivatives with respect to arbitrary functions of expected observables (Figure 4A) as follows. Consider a set of observables $\overrightarrow{O}=\{O_{1}\text{, }O_{2}\text{, }\cdots{O_{m}}\}$ where the $i^{th}$ observable corresponds to the response of the system to the $i^{th}$ value which one will use to infer an ensemble response, e.g. $O_{i}$ corresponds to the extension of a duplex subject to the $i^{th}$ value of external force. Given all $m$ values, we compute the summary statistic of interest $S$ via an arbitrary line-fitting procedure $\ell$, i.e. $\ell(\overrightarrow{O})=S$. For control parameters $\theta$, we compute $\nabla_{\theta}S$ by computing an implicit derivative for the line fitting procedure $\ell$ (see Supporting Information). Figure 4A gives an overview of our procedure, which uses implicit differentiation [62] on the fitting parameters combined with trajectory reweighting, so the entire calculation of mechanical properties becomes end-to-end differentiable. Trajectory reweighting naturally accommodates ensembles involving external forces and torques since their corresponding effect on the probability of a microstate does not depend explicitly on $\theta$ (see Supporting Information).
As a specific example, under low tension, a long molecule of dsDNA behaves like an extensible wormlike chain (WLC) [63]. In this case, both the persistence length $L_{ps}$ and the extensional modulus $S$ are extracted from fits to the WLC expression of Odijk:
with
$y=\left(\frac{FL_{0}^{2}}{L_{ps}k_{\textup{B}}\textup{T}}\right)^{1/2}.$ (2)
Here, $x$ is the extension of the duplex under force $F$; $L_{0}$ is the contour length of the duplex and $k_{\textup{B}}\textup{T}$ is the thermal energy at temperature T [64], with $k_{B}$ Boltzmann’s constant. Thus, given data for a force-extension curve calculated via a set of $m$ simulations at various forces (i.e. a set of average extensions at $m$ different forces), we compute $L_{ps}$, $S$, and $L_{0}$ using an unconstrained three parameter fit to Equation 1. Alternatively, $L_{ps}$ can be calculated passively via fitting a line to correlations in alignment as a function of separation [65]. Other physical quantities can be treated in a similar fashion: the torsional modulus $C$, the effective stretch modulus $S_{\text{eff}}$, and the twist-stretch coupling $g$ require fitting the linear dependence of changes in twist and extension as a function of applied force and torque with analytical formulas derived from the equipartition theorem [24, 66]. Finally, while structural properties are often interpreted as some mean observable of a DNA molecule, the distribution of the same observable can be viewed as a mechanical property [67]; arbitrary probability distributions can be fit to measured distributions and the calculation of summary statistics can similarly be made end-to-end differentiable. See the Supporting Information for details on each of these calculations.
@@
We first apply this method to optimize for $L_{ps}$ and $S$. We find that the most robust fitting procedure is the passive calculation described above (Figure 4B). Although the WLC fit by itself is highly sensitive to details that make optimization more challenging to use directly, we can reduce the variance of the WLC fit and successfully optimize $S$ by constraining $L_{ps}$ to the value computed via passive simulations (Figure S5).
Sensitivity analysis again reveals a strong dependence of the persistence length on the cross-stacking interaction; this differs from previous work that has focused on hydrogen-bonding and stacking interactions when fitting new persistence length data [37].
We also optimize the potential for the torsional modulus $C$ and twist-stretch coupling $g$ (Figure 4C). Strikingly, our method directly finds a value of $g$ within experimental uncertainty, requiring the model to capture the transition to overwinding behavior of dsDNA under low force (Figure 4C.II). This experimental observation previously motivated an entirely new model of DNA [24]. Our fits to $C$ and $S_{\text{eff}}$ are also within experimental uncertainty(Figure 4C.III). Again, sensitivity analysis reveals a strong dependence of all three quantities on the cross-stacking interaction (Figure S6).
Our method can be further adapted to accommodate additional mechanical properties not explicitly considered in prior oxDNA parameterizations, including those that require jointly optimizing several properties simultaneously. For example, we optimize for a distribution of propeller twists that maintains a mean value of $21.7^{\circ}$ while increasing the variance (Figure S7).
In all optimizations described in this section, we fit the sequence-averaged parameters of oxDNA; historically, sequence-specificity is added to oxDNA by separately fitting sequence-specific parameters for the stacking and hydrogen-bonding interactions as multiplicative factors. We additionally fit these parameters to an arbitrarily chosen sequence-specific persistence length (Figure S8), highlighting the flexibility of our method across sequence-averaged and sequence-specific considerations.
@@ 2.4 Thermodynamic Optimization
The third class of data to which oxDNA is fit is thermodynamic data, i.e. duplex and hairpin melting temperatures. This presents a host of challenges for traditional differentiable MD beyond those discussed so far. First, melting temperatures in oxDNA usually require fitting melting curves to expected values of discrete order parameters (i.e. number of base pairs), manifestly not differentiable via the pathwise estimator. Second, melting temperature calculations require enhanced sampling methods such umbrella sampling, in order to sample sufficient statistics of both bound and unbound structures. and this typically involves non-differentiable cluster-move Monte Carlo schemes. Third, the unbiased statistics sampled from one biased simulation at a given temperature are then used to extrapolate the corresponding statistics at a range of temperatures, from which a melting curve is finally fit.
We adapt the DiffTRE formalism to accommodate such complex calculations. In umbrella sampling, states $s_{i}$ are sampled from a biased Markov chain in which states are sampled with unnormalized probability given by
@@
where the weights $\omega_{i}$ are chosen to promote configurations along some order parameter(s) that would otherwise be inaccessible to an unweighted Monte Carlo sampling. Similarly, a simulation at temperature $T$ can be treated as a biased sample of the ensemble at a different temperature $T^{\prime}$ and the expected value of a target observable $O$ at temperature $T^{\prime}$ can be computed as
$\langle O(s)\rangle_{T^{\prime}}=\langle O(s)\exp(U(s,T)/kT-U(s,T^{\prime})/kT^{\prime})\rangle_{T}$ (4)
Note that this is a special case of Equation 3 where $\omega_{i}=\exp(U(s_{i},T)/kT)$. By reformulating the unnormalized probability of a state via Equations 3 and 4, we adapt DiffTRE to differentiate the expected number of base pairs at a given temperature $T^{\prime}$. In practice, we compute the expected ratio of bound to unbound states at each temperature $T^{\prime}$ (see Supporting Information). We can then combine these expected ratios with implicit differentiation or simple piecewise interpolation to compute a melting curve and extract $T_{m}$, thus enabling an end-to-end differentiable calculation of $T_{m}$. See Figure 5A for an overview.
We successfully apply our method to optimize melting temperatures of both DNA duplexes and hairpins. Since oxDNA matches Santa-Lucia duplex melting temperatures with high accuracy, we optimize for an arbitrarily different value of $T_{m}$ for an 8 base-pair duplex (Figure 5B.I). In principle, we can fit any (continuous and differentiable) function of the melting curve. To demonstrate this flexibility, we also fit to an arbitrarily chosen width of the melting curve for the 8 base-pair duplex (Figure 5B.II). Unlike for duplexes, oxDNA underestimates the melting temperatures of hairpins and we therefore fine-tune the parameters to match the melting temperature of a hairpin with a 6 base-pair stem with a loop comprised of 6 unpaired nucleotides as predicted by the Santa-Lucia model (Figure 5C).
Sensitivity analyses for the duplex and hairpin $T_{m}$’s (Figure S9) reveals a pronounced dependence on the cross-stacking interaction; the two most important parameters have gradients 3-4x larger than the next most important parameters. While cross-stacking is generally understood to stabilize duplexes and therefore affect melting, the relative contribution of cross-stacking to hydrogen-bonding and stacking parameters in the model is surprising. Taken together with the sensitivity analyses for structural and mechanical properties, these results suggest that the cross-stacking interaction is far more influential than previously understood in the parameterization of oxDNA.
Figure 6: Jointly fitting to multiple target properties simultaneously. A. Fitting to $S_{\text{eff}}$, pitch, and the $T_{m}$ of a hairpin with a 6 nucleotide loop and 6 base-pair stem. The target pitch is $10.25\pm 0.25$ bp/turn, the target $S_{\text{eff}}$ is $1045\pm 92$ pN, and the target $T_{m}$ is $334.5\pm 0.5$ K. B. Fitting to $S_{\text{eff}}$, $C$, pitch, and $L_{ps}$ simultaneously. The target $C$ is $436\pm 16$ $\text{pN}\cdot\text{nm}^{2}$ and the target $L_{ps}$ is $42\pm 2$ nm.
@@ 2.5 Joint Optimization
Up until now, we have considering fits to parameter on individual target properties,either structural, mechanical, or thermodynamic. Ideally, fits to oxDNA parameters jointly fit all parameters, allowing the discovery of parameters that fit all experimental properties. To carry this out, we define the objective function as a simple unweighted linear sum of RMSEs for individual properties, though our framework naturally accommodates more sophisticated objective functions (see Supporting Information).
@@
We first apply this to jointly fit $S_{\text{eff}}$, pitch, and the $T_{m}$ for the hairpin described above. This is motivated by oxDNA’s current overestimate of $S_{\text{eff}}$ and the difficulty of maintaining structural and thermodynamic properties given strong parameter interdependence. Figure 6A shows the progression of each property throughout the optimization. The optimization begins by decreasing $S_{\text{eff}}$ within experimental certainty (at the cost of the hairpin $T_{m}$) and subsequently recovers the hairpin $T_{m}$ while maintaining $S_{\text{eff}}$ and pitch within the specified uncertainty.
Second, we jointly fit $S_{\text{eff}}$, $C$, $L_{p}$, and pitch. This is motivated by difficulty in the original parameterization to capture $S$ without sacrificing accuracy in other mechanical properties. Figure 6B depicts the progression of each property throughout the optimization, and again we achieve a parameter set that models all properties within the specified uncertainty. Like Figure 6A, the optimization progresses by focusing on a single property at a time (e.g. $S_{\text{eff}}$) at the cost of unoptimized properties (e.g. $L_{ps}$), and then maintaining the optimized properties while proceeding to the next target.
@@ 3 Discussion
Differentiable programming is an emerging paradigm in scientific computing that holds great promise for bringing the successes of machine-learning to physics-based modelling. However, current methods are limited to relatively simple systems and practitioners have not yet converged on best practices for fitting complex, experimentally-relevant models. In this work, we demonstrate how state-of-the-art methods in differentiable programming can be combined with highly complex simulation routines to fit force field parameters for a well-accepted model of DNA. We successfully optimize structural, mechanical, and thermodynamic properties of DNA. Accomplishing this requires differentiating through simulation methods such as enhanced sampling, external forces and torques, and cluster-move Monte Carlo, as well as downstream calculations such as line-fitting procedures. These methods scale to long simulations of large systems, are both memory-efficient and numerically stable, directly make use of existing molecular simulation software, and permit the joint optimization of several target properties simultaneously.
@@
There is a host of existing numerical methods for fitting coarse-grained force field parameters. For example, methods like force matching, relative entropy minimization, and Boltzmann inversion are often cast as variational problems for deriving coarse-grained potentials. However, each of these methods is limited in its generality. For example, force matching is sensitive to nonlinear dependencies on the force field parameters and therefore the functional form is often restricted to spline interpolation [68]. Similarly, methods like Boltzmann inversion and relative entropy (RE) minimization are effective for matching a ground truth distribution (often derived from atomistic simulations), but the reliance on Lagrange multipliers to constrain ensemble averages (specifically in the case of RE methods) does not accommodate frustrated optimization problems (i.e. those with mutually incompatible data) [69]. In addition, the existing suite of methods cannot be naturally composed, necessitating independent parameter optimizations depending on the applicable method for a given target property. Alternatively, gradient-based methods naturally compose and accommodate frustrated systems. Moreover, gradient-based methods can immediately be integrated with the rich suite of methods developed by the machine learning community for navigating complex optimization landscapes (e.g. momentum-based optimizers, overparameterization via neural networks).
@@
There is also existing work on scaling differentiable MD to complex, experimentally-relevant force fields. Greener uses differentiable simulation with 5 ns trajectories to develop an improved force field for disordered proteins that better reproduces experimentally-measured secondary structure and radius of gyration [49], employing gradient clipping to mitigate the numerical instabilities observed in Figure 2C. Yet, in this formulation, each epoch required one day of compute on 12 GPUs, likely due to the large time cost of differentiating unrolled trajectories (Figure 2D), restricting the optimized parameter set ito only five epochs of training. The methods outlined in the present work avoid these difficulties: all optimizations presented in this work were performed on standard high performance computing CPU nodes, and gradient updates take, at most, the duration of the longest forward simulation. This enhanced efficiency permits tens to hundreds of training epochs within several hours or days at low cost.
@@
There are several key limitations of our approach. Foremost, DiffTRE is restricted to ensembles for which unnormalized probabilities of states are known and therefore we only fit to target equilibrium properties. In future work, we plan to exploit fundamental results in non-equilibrium physics (e.g. the Jarzynski equality) to differentiate relevant dynamic properties such as response to external work protocols. In the spirit of actor-critic algorithms in reinforcement learning, we also envision a more general probabilistic treatment of out-of-equilibrium properties in which the attractor of the dynamical system is parameterized and jointly inferred. There is also great promise in approximate methods of gradient estimation that may enable tractable gradient calculation through unrolled trajectories [70]. A second limitation of our work arises when jointly optimizing for several target properties. In principle, gradients of individual loss terms can conflict, trapping the optimization procedure in local minima. A promising direction to alleviate this limitation is the method introduced by Liu et al. for computing conflict-free update directions given multiple loss terms [71].
@@
There are many additional opportunities for future work. For example, we recently introduced a method for directly designing biopolymers via simulation [72] and our JAX implementation of oxDNA is amenable to this method. Second, existing coarse-grained models are often developed to reproduce models that are themselves approximations, such as the nearest-neighbor model and atomistic simulations. In our framework, one could fit a model of DNA or RNA directly to experimental data (e.g. the optical melting experiments underlying the nearest-neighbor model) rather than the model itself. Finally, we are excited to work with experts in RNA and DNA modelling to fit new coarse-grained models incorporating full datasets of structural, mechanical, and thermodynamic data. Such efforts will benefit from a standardized library of target behavior and corresponding simulation protocols for direct integration with our optimization framework.

### v2

@@ 2.3 Mechanical Optimization
To solve these challenges, we extend DiffTRE to compute derivatives with respect to arbitrary functions of expected observables (Figure 3A) as follows. Consider a set of observables $\overrightarrow{O}=\{O^{1}\text{, }O^{2}\text{, }\ldots\text{, }{O^{m}}\}$ where the $i^{th}$ observable corresponds to the response of the system to the $i^{th}$ value used to infer an ensemble response, e.g. $O^{i}$ corresponds to the extension of a duplex subject to the $i^{th}$ value of external force. Given all $m$ values, we compute the summary statistic of interest $S$ via an arbitrary line-fitting procedure $\ell$, i.e. $\ell(\overrightarrow{O})=S$. Figure 3A gives an overview of our procedure for computing $\nabla_{\theta}S$ for control parameters $\theta$. We use implicit differentiation [68] on the fitting parameters and combine this with trajectory reweighting, rendering the entire calculation of mechanical properties end-to-end differentiable. Details of our implementation can be found in the Supporting Information. Trajectory reweighting naturally accommodates ensembles involving constant external forces and torques since their corresponding effect on the probability of a microstate does not depend explicitly on $\theta$ (see Equation S63).
Figure 4: Fitting to a target propeller twist distribution. A. We optimize the parameters to achieve a target distribution of propeller twists. Here, we consider the case of increasing the variance ($\sigma^{2}=5.5$) while maintaining the mean ($\mu=21.7^{\circ}$). The loss is measured as the KL-divergence between the simulated distribution and the target distribution. The initial and target distributions are represented as Gaussians. We also plot histograms of the sampled propeller twists with the initial parameters and the optimized parameters. The initial parameters yield a distribution with $\sigma^{2}=1.5$. B. The progression of the mean, variance, and KL-divergence (i.e. the loss) throughout the optimization. Target values are depicted as dashed red lines.
As a specific example, $L_{ps}$ can be calculated passively by treating DNA as an infinitely long, semi-flexible polymer (Figure 3B-I). Under this assumption, correlations in alignment decay exponentially with separation and $L_{ps}$ obeys the following equality:
where $\mathbf{l_{0}}$ denotes the vector between the first two monomers and $\langle\mathbf{l_{m}}\cdot\mathbf{l_{0}}\rangle$ the correlation between the tangent vectors corresponding to the first and $n^{th}$ monomers [39, 73]. As shown in Figure 3B-II, our framework provides a robust means of optimizing $L_{ps}$; see the Supporting Information for simulation and calculation details. As in the structural case, sensitivity analysis also reveals a strong dependence of $L_{ps}$ on the cross-stacking interaction (Figure 3B-III); this is surprising given that previous work focused on hydrogen-bonding and stacking interactions when fitting new persistence length data [26].
We demonstrate the generality of our approach by fitting to additional mechanical properties: the torsional modulus $C$, the effective stretch modulus $S_{\text{eff}}$, and the twist-stretch coupling $g$ (Figure 3C-I). Computing these properties requires fitting the linear dependence of changes in twist and extension as a function of applied force and torque with analytical formulas derived from the equipartition theorem [20, 74]. For these optimizations, we sample reference states using the LAMMPS implementation of oxDNA [75] given existing simulation protocols [20]. Strikingly, our method yields a parameter set that successfully models the overwinding of DNA under low force, represented by a negative value of $g$, within experimental error (Figure 3C-II). The original oxDNA model incorrectly predicts underwinding. In addition, we are able to achieve $C$ and $S_{\text{eff}}$ values that agree with experiments (within experimental uncertainty); see Figure 3C-III. Again, sensitivity analysis reveals a strong dependence of all three quantities on the cross-stacking interaction (Figure S7).
@@
The framework outlined above can also be applied to fit distributions of equilibrium structural properties, as this requires fitting the functional form of a probability distribution to a histogram of sampled quantities. To illustrate this, we fit to a target equilibrium distribution of propeller twists (Figure 4). Likewise, calculating worm-like chain (WLC) properties requires fitting a non-linear force-extension response [76]. Though the WLC fit is highly sensitive to small changes in measured extensions and therefore challenging to optimize directly, we can successfully optimize the stretch modulus $S$ by constraining the fit to the value of $L_{ps}$ computed via passive simulations (Figure S6). Lastly, sequence-specificity can substantially affect response to force and torque. Sequence-specificity is modelled in oxDNA by fitting multiplicative parameters for the stacking and hydrogen-bonding interactions. We fit these parameters to an arbitrarily chosen $L_{ps}$ (Figure S8), highlighting the flexibility of our method across sequence-averaged and sequence-specific considerations.
Figure 6: Jointly fitting to multiple target properties simultaneously via conflict-free gradients. A. Conflict-free gradients provide a means of imposing hard constraints. I. Given a loss function expressed as the sum of individual loss terms, directly differentiating the total loss may yield a gradient that conflicts with the gradients of individual terms. Instead, the conflicting components of individual gradients can be projected out to yield a conflict-free gradient that is analytically guaranteed to minimize all subterms. This figure is adapted from Ref. [59]. II. A detailed schematic of a conflict-free update for an arbitrary number of hard constraints, in which one first computes the gradients corresponding to individual loss terms and then maps these individual gradients to a single conflict-free gradient via a conflict-resolution scheme. There are many choices of conflict-resolution schemes developed in the context of multi-task learning. We use the ConFIG method of Zhang et al. [59]. In B-D, all optimizations are performed with and without conflict-free updates, using identical hyperparameters. B. Fitting oxDNA parameters to $S_{\text{eff}}$, pitch, and the $T_{m}$ of a hairpin with a 6 nucleotide loop and 6 base-pair stem. The target pitch is $10.25$ bp/turn, the target $S_{\text{eff}}$ is $1045$ pN, and the target $T_{m}$ is $334.5$ K. Iterations in which reference states are resampled are represented as scatter points. Reference state resampling information is omitted from C-D. C. Fitting oxRNA parameters to $L_{ps}$, pitch, inclination, and rise. The target $L_{ps}$ is $48$ nm, the target rise is $0.28$ nm, the target pitch is $10.9$ bp/turn, and the target inclination is $15.5^{\circ}$. D. Fitting the parameters of a custom energy function for DNA-protein complexes to minimize the root mean square deviation (RMSD) between simulated structures and three experimentally-derived zinc finger structures from the Protein Data Bank (PDB). Under this energy function, DNA energetics are described by oxDNA, protein energetics are described by an anisotropic network model (ANM), and DNA-protein interactions are described by a sequence-specific Morse potential (see Materials and Methods).
@@ 2.4 Thermodynamic Optimization
The third class of data to which coarse-grained models are fit is thermodynamic data, e.g. melting temperatures. This presents a host of challenges for traditional differentiable MD beyond those discussed so far. For example, melting temperatures usually require fitting melting curves to expected values of discrete order parameters (e.g. number of base pairs), manifestly not differentiable via the pathwise estimator. Additionally, melting temperature calculations typically require enhanced sampling methods such as umbrella sampling given the rare nature of the binding/unbinding events and the necessity of sampling multiple such transitions in order to amass sufficient statistics. Methods for performing such biased simulations historically involve non-differentiable Monte Carlo schemes, e.g. oxDNA applies a bespoke cluster-move Monte Carlo algorithm [34]. Lastly, statistics sampled at a given temperature are typically extrapolated to derive statistics at a range of temperatures, necessitating similar unbiasing principles.
To overcome these challenges, we extend the DiffTRE formalism to accommodate such complex calculations. In umbrella sampling, states $s_{i}$ are sampled from a biased Markov chain with unnormalized probability given by
@@
where the weights $\omega_{i}$ are chosen to promote configurations along some order parameter(s) that would otherwise be virtually inaccessible to an unweighted Monte Carlo sampling. Similarly, a simulation at temperature $T$ can be treated as a biased sample of the ensemble at a different temperature $T^{\prime}$ and the expected value of a target observable $O$ at temperature $T^{\prime}$ can be computed as
$\langle O(s)\rangle_{T^{\prime}}=\langle O(s)\exp\left(U(s,T)/kT-U(s,T^{\prime})/kT^{\prime}\right)\rangle_{T}$ (3)
Note that this is a special case of Equation 2 where $\omega_{i}=\exp(U(s_{i},T)/kT)$. Summary statistics of the temperature response may then be computed by line-fitting procedures and differentiated via implicit differentiation, as with mechanical properties. See Figure 5A for an overview of this framework.
This framework enables the differentiable calculation of the melting temperature $T_{m}$ of duplexes and hairpins in oxDNA. This is achieved by calculating the expected ratio of bound to unbound states at a range of temperatures and computing a melting curve to extract $T_{m}$ (see Supporting Information). We successfully apply our method to optimize melting temperatures of both DNA duplexes and hairpins. Since oxDNA matches experimental duplex melting temperatures with high accuracy [26, 69], we optimize for an arbitrarily different value of $T_{m}$ for an 8 base-pair duplex (Figure 5B-I). In principle, we can fit any (continuous and differentiable) function of the melting curve. To demonstrate this flexibility, we also fit to an arbitrarily chosen width of the melting curve for the 8 base-pair duplex (Figure 5B-II). Unlike for duplexes, oxDNA underestimates the melting temperatures of hairpins. We therefore fine-tune the parameters to match the melting temperature of a hairpin with a 6 base-pair stem with a loop comprised of 6 unpaired nucleotides as predicted by the Santa-Lucia model (Figure 5C) [77].
Sensitivity analysis once more reveals a pronounced dependence on the cross-stacking interaction (Figure S9). While cross-stacking is generally understood to stabilize duplexes and therefore affect melting, the relative contribution of cross-stacking compared to hydrogen-bonding and stacking in the model is surprising. Taken together with the sensitivity analyses for structural and mechanical properties, these results suggest that the cross-stacking interaction is far more influential than previously understood in the parameterization of oxDNA.
@@ 2.5 Joint Optimization
Until now, we have considered fitting parameters to individual target properties – either structural, mechanical, or thermodynamic. Ideally, parameters should be simultaneously optimized using a joint loss function that describes multiple targets. To carry this out, we first define the objective as the simple linear sum of root mean squared errors (RMSEs) for individual properties, though our framework naturally accommodates more sophisticated objective functions (see Supporting Information). However, this vanilla formulation has the property that gradients of one target property may conflict with another, resulting in the improvement of one term at the cost of another.
To address this, we turn to conflict-free gradient methods from multi-task learning problems in machine learning. These methods compute an update direction by projecting out conflicting components of individual gradients, ensuring that multiple loss terms can be minimized simultaneously without interference (Figure 6A). We use the conflict-resolution method of Liu et al. [59], though there are many alternatives [55, 56, 78, 79]. In the context of fitting coarse-grained potentials, this allows us to define a notion of hard constraints by which we ensure that each term included in the conflict-free update decreases monotonically (up to momentum artifacts in the numerical optimizer). See Materials and Methods for details on constructing a conflict-free update.
@@
To illustrate this approach, we jointly fit $S_{\text{eff}}$, pitch, and the $T_{m}$ for the hairpin described above. This is motivated by oxDNA’s current overestimate of $S_{\text{eff}}$ and the difficulty of maintaining structural and thermodynamic properties given strong parameter interdependence. We show the progression of each property throughout the optimization in Figure 6B. Without conflict-free updates, the optimization begins by decreasing $S_{\text{eff}}$ and pitch at the cost of the hairpin $T_{m}$, subsequently recovering the hairpin $T_{m}$ while maintaining $S_{\text{eff}}$. Conflict-free updates make no such sacrifices, converging to the target values in roughly half as many gradient updates. Second, we jointly fit $S_{\text{eff}}$, $C$, $L_{ps}$, and pitch (Figure S10). This is motivated by difficulty in the original parameterization of capturing $S$ without sacrificing accuracy in other mechanical properties. Again, we find that conflict-free updates yield a more stable and efficient means of optimization.
2.6 Extension to other models
We have focused so far on reparameterizing the oxDNA model as a representative case study of our framework, but anticipate our approach can be extended to a broad family of coarse-grained biomolecular models. To illustrate this, we conclude by treating two other example systems: RNA (Figure 6C), a notoriously challenging subject of molecular modeling [80, 81, 82, 83, 84, 85, 86], and DNA-protein complexes (Figure 6D), a subject of broad therapeutic interest [87, 88, 89]. These systems highlight the utility of conflict-free gradients beyond merely accelerating optimization.
For RNA, we consider the oxRNA model, a sister model to oxDNA that captures its structure and duplex thermodynamics but poorly models mechanical properties (see Materials and Methods). We optimize oxRNA to match the experimentally-derived $L_{ps}$ while maintaining key structural properties already captured in the model (i.e. rise, pitch, and inclination). All quantities are computed as in Ref. [24]. We find that conflict-free gradient updates yield significantly more stable optimization than vanilla gradients. Using direct gradients, agreement with the target pitch value is degraded while optimizing towards the target $L_{ps}$ value and, after recovering the target pitch, $L_{ps}$ oscillates around the target value, likely due to frustration with other loss terms. Conflict-free gradient updates, however, yield a stable optimization and a parameter set that matches the target $L_{ps}$ with low variance and without compromising existing agreement with structural properties, as shown in Figure 6. To the best of our knowledge, these optimized updated parameters represent the most coarse-grained RNA model to capture these four properties simultaneously [90, 91, 92, 93].
For DNA-protein complexes, we introduce a new force field that parameterizes DNA-protein interactions via a custom sequence-specific Morse potential (see Materials and Methods). This is an extension of the model introduced in Ref. [25], in which DNA and protein molecules interact solely via hard-sphere repulsion. We circumvent the need to manually derive forces by simulating in JAX-MD, underscoring the flexibility offered by our framework for experimenting with custom functional forms. We fit model parameters directly to structures deposited in the Protein Data Bank, motivated by recently developed machine-learning models that fit directly to such structures [13, 94]. We target three zinc fingers with the loss for each zinc finger defined as the average RMSD between the simulated structure and the coarse-grained representation generated by oxView [95]. Despite the optimization problem being highly overparameterized, vanilla gradients become trapped in a local minimum which improves agreement with two structures while severely degrading performance with the third. Alternatively, conflict-free gradients serve as stable updates that yield a substantially lower total loss without compromising the accuracy of a single structure (Figure 6).
@@ 3 Discussion
Differentiable programming is an emerging paradigm in scientific computing that holds great promise for bringing the successes of machine learning to physics-based modeling. However, current methods are limited to relatively simple systems and practitioners have not yet converged on best practices for fitting complex, experimentally-relevant models.
In this work, we demonstrate how state-of-the-art methods in differentiable programming can be combined with highly complex simulation routines. We successfully optimize structural, mechanical, and thermodynamic properties of DNA, exceeding the accuracy of the original parameterization for a range of targets and introducing and successfully fitting to qualitatively new objectives, such as the width of melting curves. Accomplishing this requires differentiating complex simulation methods such as enhanced sampling, use of external forces and torques, and bespoke Monte Carlo algorithms, as well as downstream calculations such as line-fitting procedures. Our approach scales to long simulations of large systems, is both memory-efficient and numerically stable, and directly makes use of existing molecular simulation software. By adapting conflict-free gradient methods from reinforcement learning, we provide a framework for systematically fitting to multiple data sources simultaneously, a principle which has heretofore been lacking in the molecular modeling community. We demonstrate the generality and flexibility of our method by performing joint optimizations for models of RNA and DNA-protein complexes.
@@
There is a host of existing numerical methods for fitting coarse-grained force field parameters. For example, methods like force matching, relative entropy minimization, and Boltzmann inversion are often cast as variational problems for deriving coarse-grained potentials. However, each of these methods is limited in its generality. For example, force matching is sensitive to nonlinear dependencies on the force field parameters and therefore the functional form is often restricted to spline interpolation [96]. Similarly, methods like Boltzmann inversion and relative entropy (RE) minimization are effective for matching a ground truth distribution (often derived from atomistic simulations), but the reliance on Lagrange multipliers to constrain ensemble averages (specifically in the case of RE methods) does not accommodate frustrated optimization problems (i.e. those with mutually incompatible data) [97]. In addition, the existing suite of methods cannot be naturally composed, necessitating independent parameter optimizations depending on the applicable method for a given target property.
Alternatively, gradient-based methods for individual target properties can be naturally composed, either (i) by combining individual objective functions into a global objective function or (ii) via more sophisticated methods like the conflict resolution scheme employed in this work. Moreover, gradient-based methods can immediately be integrated with the rich suite of tools developed by the machine learning community (e.g. momentum-based optimizers, overparameterization via neural networks) for navigating complex optimization landscapes (e.g. frustrated systems). Recent work by Fuchs et al. demonstrates the effectiveness of these tools for optimizing molecular force fields by developing an ML-inspired ecosystem for fitting neural-network (NN) potentials to biomolecular experimental and simulation data [98].
@@
There is also existing work on scaling differentiable MD to complex, experimentally-relevant force fields. Greener uses differentiable simulation with 5 ns trajectories to develop an improved force field for disordered proteins that better reproduces experimentally-measured secondary structure and radius of gyration [32], employing gradient clipping to mitigate numerical instabilities like those observed in Figure S2D. Yet, in this formulation, each epoch required one day of compute on 12 GPUs, likely due to the large time cost of differentiating unrolled trajectories (Figure S2B), restricting the optimized parameter set to only five epochs of training. The methods outlined in the present work avoid these difficulties: all optimizations presented here were performed on standard high-performance computing CPU nodes, and gradient updates take, at most, the duration of the longest forward simulation. This enhanced efficiency permits tens to hundreds of training epochs within several hours or days at low cost.
@@
There are several limitations of our approach. Foremost, DiffTRE is restricted to ensembles for which unnormalized probabilities of states are known and therefore we only fit to target equilibrium properties. In future work, we plan to exploit fundamental results in non-equilibrium physics (e.g. the Jarzynski equality [99]) to differentiate dynamical properties such as reaction rates and response to external work protocols. In the spirit of actor-critic algorithms in reinforcement learning, we also envision a more general probabilistic treatment of out-of-equilibrium properties in which the attractor of the dynamical system is parameterized and jointly inferred. There is also great promise in approximate methods of gradient estimation that may enable tractable gradient calculation through unrolled trajectories [100].
@@
There are many additional opportunities for future work. For example, we recently introduced a method for directly designing biopolymers via simulation [101] and our JAX implementations of oxDNA and oxRNA are amenable to this method. Second, existing coarse-grained models are often developed to reproduce models that are themselves approximations, such as the nearest-neighbor model and atomistic simulations. In our framework, one could fit a model of DNA or RNA directly to experimental data (e.g. the optical melting experiments underlying the nearest-neighbor model) rather than the model itself. Finally, we are excited to work with experts in biomolecular modelling to fit new coarse-grained models incorporating full datasets of structural, mechanical, and thermodynamic data. Such efforts will benefit from a standardized library of experimental target behaviors and corresponding simulation protocols for direct integration with our optimization framework.